# Ambiguous Resonances in Multipulse Quantum Sensing with Nitrogen Vacancy Centers

### v2

@@ Abstract
Dynamical decoupling multipulse sequences can be applied to solid state spins for sensing weak oscillating fields from nearby single nuclear spins. By periodically reversing the probing system’s evolution, other noises are counteracted and filtered out over the total evolution. However, the technique is subject to intricate interactions resulting in additional resonant responses, which can be misinterpreted with the actual signal intended to be measured. We experimentally characterized three of these effects present in single nitrogen vacancy centers in diamond, where we also developed a numerical simulations model without rotating wave approximations, showing robust correlation to the experimental data. Regarding centers with the 15N nitrogen isotope, we observed that a small misalignment in the bias magnetic field causes the precession of the nitrogen nuclear spin to be sensed by the electronic spin of the center. Another studied case of ambiguous resonances comes from the coupling with lattice 13C nuclei, where we reconstructed the interaction Hamiltonian based on echo modulation frequencies and used this Hamiltonian to simulate multipulse sequences. Finally, we also measured and simulated the effects from the free evolution of the quantum system during finite pulse durations. Due to the large data volume and the strong dependency of these ambiguous resonances with specific experimental parameters, we provide a simulations dataset with a user-friendly graphical interface, where users can compare simulations with their own experimental data for spectral disambiguation. Although focused with nitrogen vacancy centers and dynamical decoupling sequences, these results and the developed model can potentially be applied to other solid state spins and quantum sensing techniques.
@@ I Introduction
The second quantum revolution is driving sensing technology [1] to ever-increasing precision and accuracy, thus opening doors to novel measurement possibilities [2, 3, 4, 5], reaching the detection of single nuclear spins at room temperature [6, 7, 8, 9]. This quantum sensing advantage relies on the fact that quantum properties are extremely sensitive to their environment, which contrarily is a main obstacle for quantum computing — the issue of decoherence [10]. However, such key element is also a vulnerability, as quantum sensors are susceptible to complex interactions with their environment, leading to responses that can be misinterpreted with the actual signal intended to be measured.
In this study, we consider these ambiguous resonant responses which affect multipulse dynamical decoupling (DD) pulse sequences. The technique consists in applying repeated periodic reversals of the quantum evolution of the system, such that dephasings from the environment are counteracted over the entire time evolution. The simplest possible decoupling sequence, known as Hahn echo [11], consists of a single $\pi$-pulse between two free-evolutions with same duration [$\tau-\pi-\tau$], being originally developed as a way to extend coherence of spin ensembles in nuclear magnetic resonance (NMR). The Hahn echo decay signal can already be used to sense coupled spins [12], however as the number of pulses and free evolution is extended (CPMG sequence [13]), the sequence acts as a narrow frequency or time period filter which cancels out AC signals, except for the frequency corresponding to double of the pulse separation $f_{0}=1/(2\tau)$ and odd fractions of it [14, 15, 16]: $f_{0}/3$, $f_{0}/5$, $f_{0}/7$… Beyond the Hahn echo and CPMG sequences, a further improvement can be made by intercalating $\pi$-pulses over the $x$ and $y$ axes to decouple noises acting on the two different axes. One of these is the XY8-$M$ sequence, with eight $\pi_{x,y}$-pulses repeated $M$ times as schematically represented in Fig. 1. By decoupling noises acting on different axes, weak signals from single nuclear spins can be filtered from much stronger background fields [6, 7, 8], representing a remarkable improvement compared to conventional NMR, which requires large nuclear ensembles and much stronger bias magnetic fields. This improvement come at the expense that the probing system is now intrinsically non-classical, hence having knowledge of its time evolution is indispensable for the resonant signal analysis. To overcome these challenges, an in-depth characterization of such ambiguous resonances experienced by the quantum sensor is necessary for the continued development of the field.
Figure 1: The XY8-$M$ dynamical decoupling sequence is composed of eight alternated $\pi_{x,y}$-pulses repeated $M$ times, which counteracts oscillating magnetic field noises except for frequencies corresponding to half of the pulse separation $f_{0}=1/(2\tau)$ and odd fractions of it. This way, it can be applied in a probing spin to sense weak oscillating fields from single nuclear spins.
@@
Figure 2: (a) The NV center is composed of a nitrogen substitution adjacent to a vacancy in the carbon lattice of diamonds. In general, the external magnetic field $\vec{B}_{0}$ forms an angle $\theta_{0}$ with the NV axis. 13C isotopes with spin 1/2 can also be present near the center. (b) Energy level scheme of NV-. The electronic spin can be initialized to $m_{s}=0$ by optical pumping with green laser, while the state population can be measured by the resulting red fluorescence. Coherent control between $m_{s}=0\leftrightarrow\pm 1$ states is achieved with resonant microwave pulses. (c) Schematic representation of the experimental setup. An air objective focuses the green laser on a single NV, which is also used to collect its red fluorescence. The bias magnetic field $\vec{B}_{0}$ is applied with a permanent magnet, while microwave pulses and radio signals are transmitted through a wire on top of the diamond.
@@
Quantum sensing DD techniques could in principle be applied to any two-level quantum system. Nonetheless, they excel with color center defects in solids [17], given their versatility to sense different physical quantities and relative ease of experimental realization. Among them, the nitrogen vacancy (NV) center in diamonds has been an epicenter of research due to its remarkable properties at a large range of temperatures and compatibility with biological systems [18, 19], unlike for instance, superconducting quantum interference devices (SQUIDs) [20]. An NV center is a substitutional defect in the diamond lattice where a carbon atom is replaced by nitrogen 14N ($I^{n}=1$) or 15N ($I^{n}=1/2$) adjacent to a vacancy [Fig. 2 (a)]. The center can accept an extra electron from the lattice, acquiring a negative charge state NV- and forming a spin triplet ($S=1$) or singlet ($S=0$) configuration with the unpaired valence electron of the nitrogen, resulting in well-defined quantum states within the bandgap of diamond [Fig. 2 (b)]. Since our work is concerned with NV- centers, in the following we denote them as NV for simplicity.
The ground triplet states 3A can be driven to the excited 3E states with a non-resonant green laser excitation via the phonon side band. Which then leads to a polarization of the electronic spin state $m_{s}$ by an optical pumping mechanism via the singlet metastable states, with a preferential decay to $m_{s}=0$ [21]. Based on the same mechanism, the resulting fluorescence can be used to resolve the population difference between $m_{s}=0$ and $m_{s}=\pm 1$, since the latter fluoresce less intensively. Finally, transitions between $m_{s}=0$ and $m_{s}=\pm 1$ levels of the ground triplet states 3A can be coherently driven by resonant microwaves (MW) with controllable duration and phase [22]. This way, a DD sequence consists of three parts: laser initialization, MW manipulation and fluorescence readout. For NVs, $\pi$/2-pulses are also included before and after DD sequences, in order to drive the electronic spin to and from the plane perpendicular to the quantization axis.
@@
In the context of single NV multipulse sequences, ambiguous resonances can arise from different factors. We begin with Sec. II giving a general theoretical framework used in our simulations, following up with a description of the experimental and simulation methods in Sec. III. We then discuss each ambiguous resonance: external magnetic field misalignment and hyperfine interaction with 15N nuclei (Sec. IV), coupling with lattice 13C having $I^{c}=1/2$ (Sec. V) and free evolution of the system during finite pulses duration [16] (Sec. VI). In sight of the numerous possible ambiguous resonances specific to experimental parameters, we conclude with Sec. VII providing a dataset with graphical interface where simulations can be visualized and compared with the user’s experimental data for spectral disambiguation of resonances. Although the discussion is focused on NVs, some of these effects are also relevant to other color centers and quantum systems in general.
@@ II Theoretical Framework
Differently from conventional NMR, in DD quantum sensing, the probing system is an inseparable part from the target spin and the exact Hamiltonian of the quantum sensor needs to be taken into account. The total Hamiltonian of a NV in a multipulse sequence can decomposed in three parts as $H=H_{0}+H_{1}(t)+H_{2}(t)$, where $H_{0}$ is the internal component, $H_{1}(t)$ is the control Hamltonian describing the interaction with the MW pulses and $H_{2}(t)$ the sensing interaction itself. For the first internal component, we consider a model of the ground state triplet, already extensively studied and characterized [23, 24, 25]. It can be expressed in the Hilbert space of the electronic and nitrogen nuclear spins $\mathcal{H}_{S}\otimes\mathcal{H}_{I^{n}}$ in factors of decreasing energy as
$H_{0}=D\left({S_{z}}^{2}-\frac{S^{2}}{3}\right)+E\left({S_{x}}^{2}-{S_{y}}^{2}\right)$ (1.1)
$+\gamma^{e}\vec{B}_{0}\cdot\vec{S}$ (1.2)
@@
where the $z$ axis is taken along the NV axis. The first line (Eq. 1.1) corresponds to the zero-field splitting due to the dipolar interaction between the two electrons, where $D=2.87$ GHz is the axial component. The parameter $E$ is the off-axis component and is typically negligible, except in nanodiamonds where the crystal strain is significant [25]. The second and fourth lines (Eqs. 1.2 and 1.4) correspond to Zeeman interactions with the external magnetic field $\vec{B_{0}}$ for the electronic spin and nitrogen nuclear spin respectively, with gyromagnetic ratios $\gamma^{e}=28025$ MHz/T, $\gamma^{n}=-4.316$ MHz/T for 15N and $\gamma^{n}=3.077$ MHz/T for 14N. Here we consider a general case where $\vec{B_{0}}$ is not necessarily aligned with the NV axis, as usually assumed. The third line (Eq. 1.3) is the hyperfine interaction between spins which is diagonal in this symmetry referential, with $a_{\perp}^{n}=-2.70$ MHz and $a_{\parallel}^{n}=-2.14$ MHz for 14N, while $a_{\perp}^{n}=3.65$ MHz and $a_{\parallel}^{n}=3.03$ MHz for 15N. When considering 14N, there is an extra quadrupole term given by $Q(I_{z}^{n})^{2}$, where $Q=-5.01$ MHz. Although small, the presence of the quadrupole interaction fundamentally changes the dynamics of the system depending on the nitrogen isotope, as the interaction tends to fix the nuclear spin orientation and suppress the Larmor precession of 14N nuclear spins for magnetic fields $B_{0}\ll Q/\gamma^{n}=1.63$ T [12].
Considering the control Hamiltonian $H_{1}(t)$, for an oscillating field along a vector $\vec{B_{1}}$, with angular frequency $\Omega_{MW}$ and phase $\phi_{1}$, it can be described as
@@
Typically, a rotating wave approximation (RWA) is assumed [26]. Thus, by setting the pulse frequency in resonance with one of the $m_{s}=0\leftrightarrow\pm 1$ transitions, $H_{1}(t)$ can be treated as a time independent operator which induces a rotation on the Bloch sphere of the respective manifold, where the pulse duration controls the angle and $\phi_{1}$ the axis of rotation in the rotating frame. Although this approximation often results in accurate descriptions of the electronic states transitions, it can overlook important interactions which lead to the ambiguous resonant responses. In particular, the $S_{x,y}$ components of the electronic Zeeman and the perpendicular terms of the hyperfine interactions $a_{\perp}^{n}\left(S_{x}I_{x}^{n}+S_{y}I_{y}^{n}\right)$ do not commute with the rotation operator and result in extra resonance peaks (Sec. IV). This time dependency resulting from the non-commuting terms could still be handled as a perturbation to the secular part of the Hamiltonian up to 2nd order of the dominant term $(1/D)^{p}$ [12, 27, 28]. However, this approach is also limited, as it overlooks the possibility of double quantum transitions caused by multi-frequency or broad band excitations [29], as well as implies small bias fields $|\gamma_{e}\vec{B}_{0}|\ll D$ and weak driving regimes $\gamma_{e}B_{1}\ll D$, which is not always experimentally desired due to extended pulse durations [30]. Furthermore, perturbation theory, in such circumstances, restricts on calculating approximated eigenvalues for the perturbed Hamiltonian, ignoring exact characteristics of the MW field and the free-evolution of the system during finite pulse durations. Which also leads to other ambiguous resonances (Sec. VI) [16].
@@
For a single nuclear spin in the strong coupling regime but without contact term [9], the sensing Hamiltonian $H_{2}$ can be simply expressed in an extended Hilbert space to include the sensed target spin $\mathcal{H}_{S}\otimes\mathcal{H}_{I^{n}}\otimes\mathcal{H}_{I^{t}}$. For $I^{t}=1/2$ and neglecting the interaction between the nitrogen and target nuclear spin, this Hamiltonian is given by a hyperfine coupling between the electronic spin operator and the nuclear spin $\vec{I}^{t}$ added to the Zeeman interaction with the external magnetic field
$H_{2}=\vec{S}\cdot A^{t}\cdot\vec{I}^{t}-\gamma^{t}\vec{B}_{0}\cdot\vec{I}^{t}.$ (3)
@@
Diagonalizing the total Hamiltonian $H_{0}+H_{2}$ results in distinct target nuclear spin Larmor frequencies $\tilde{\omega}_{m_{s}}^{t}$, dependent on the electronic spin state. Given that the hyperfine matrix $A^{t}$ must be symmetrical, the sensing interaction is completely described by its six components plus the gyromagnetic ratio $\gamma^{t}$. The same Hamiltonian can also be used to describe an ensemble of nuclear spins, provided that their states are represented by density operators. Apart from nuclear spins, it is also desirable to sense oscillating classical fields. Particularly for this study, an externally applied RF signal has the advantage of having controllable intensity and frequency. In this case, the interaction is described in terms of a field $\vec{B}_{2}(t)$ with frequency $f_{0}$ as
@@
In a simplified manner, an ensemble of nuclear spins could also be thought of in terms of a semi-classical oscillating field [9, 31]. Despite their average magnetization being close to zero at room temperature, their collective precession around $\vec{B}_{0}$ results in an effective oscillating magnetic field with statically varying amplitude and phase, but with frequencies centered around the effective nuclear Larmor frequencies $\tilde{\omega}_{m_{s}}^{t}$ at each electronic state $m_{s}$. On the other hand, a weak classical RF signal could emulate a precessing nuclear spin and be approximately described in terms of an effective hyperfine interaction. Hence, an analogy can be drawn between the quantum and semi-classical descriptions from Eqs. 3 and 4, where the interaction is either represented in terms of the six hyperfine components plus gyromagnetic ratio $\{A^{c}_{ij},\gamma^{s}\}$ or the three nuclear precision frequencies plus four parameters of the effective field vector $\{\tilde{\omega}_{m_{s}}^{t},\phi_{2},\vec{B}_{2}\}$, respectively. Strictly speaking, both Hamiltonians describe fundamentally different processes, yet they share mathematical similarities and can lead to resembling effects at low order XY8-$M$ sequences, as further shown in the Supplementary Materials Sec. S5.

### v3

@@ Abstract
Dynamical decoupling multipulse sequences can be applied to solid state spins for sensing weak oscillating fields from nearby single nuclear spins. By periodically reversing the probing system’s evolution, other noises are counteracted and filtered out over the total evolution. However, the technique is subject to intricate interactions resulting in additional resonant responses, which can be misinterpreted with the actual signal intended to be measured. We experimentally characterized three of these effects present in single nitrogen vacancy centers in diamond, where we also developed a numerical simulations model without rotating wave approximation, showing robust correlation to the experimental data. Regarding centers with the 15N nitrogen isotope, we observed that a small misalignment in the bias magnetic field causes the precession of the nitrogen nuclear spin to be sensed by the electronic spin of the center. Another studied case of ambiguous resonances comes from the coupling with lattice 13C nuclei, where we used the echo modulation frequencies to obtain the interaction Hamiltonian and then utilized the latter to simulate multipulse sequences. Finally, we also measured and simulated the effects from the free evolution of the quantum system during finite pulse durations. Due to the large data volume and the strong dependency of these ambiguous resonances with specific experimental parameters, we provide a simulations dataset with a user-friendly graphical interface, where users can compare simulations with their own experimental data for spectral disambiguation. Although focused with nitrogen vacancy centers and dynamical decoupling sequences, these results and the developed model can potentially be applied to other solid state spins and quantum sensing techniques.
@@ I Introduction
The second quantum revolution is driving sensing technology [1] to ever-increasing precision and accuracy, thus opening doors to novel measurement possibilities [2, 3, 4, 5, 6, 7, 8, 9]. This quantum sensing advantage relies on the fact that quantum properties are extremely sensitive to their environment. However, such key element is also a vulnerability, as quantum sensors are susceptible to complex interactions with their environment, leading to responses that can be misinterpreted with the actual signal intended to be measured.
In this study, we consider these ambiguous resonant responses which affect multipulse dynamical decoupling (DD) pulse sequences. More specifically, the XY8-$M$ [10, 11] and other similar sequences, as schematically represented in Fig. 1 (a).
@@
Figure 1: (a) The XY8-$M$ dynamical decoupling sequence is composed of eight alternated $\pi_{x,y}$-pulses repeated $M$ times, which counteracts oscillating magnetic field noises except for frequencies corresponding to half of the pulse separation $f_{0}=1/(2\tau)$ and odd multiples of it. This way, it can be applied with a probing spin to sense weak oscillating fields from single nuclear spins. (b) The NV center is composed of a nitrogen substitution adjacent to a vacancy in the carbon lattice of diamonds. This results in an electronic spin triplet or singlet configurations, with well-defined quantum states within the band-gap of diamond. In general, the external magnetic field $\vec{B}_{0}$ forms an angle $\theta_{0}$ with the NV axis. 13C isotopes with spin 1/2 can also be present near the center. (c) Schematic representation of the experimental setup. An air objective focuses the green laser to polarize the NV in the $m_{s}=0$ state, which is also used to collect its red fluorescence and determine the $m_{s}$ state. The bias magnetic field $\vec{B}_{0}$ is applied with a permanent magnet, while microwave pulses for coherent control of the electronic spin are transmitted through a wire on top of the diamond.
By periodically reversing the quantum evolution of the system with intercalated $\pi_{x}$ and $\pi_{y}$ pulses, dephasings from the environment are counteracted over the total evolution. Thus, making the sequence into a narrow filter for frequencies corresponding to double of the pulse separation $f_{0}=1/(2\tau)$ and odd multiples of it [12, 13]: $3f_{0}$, $5f_{0}$, $7f_{0}$. Like so, weak signals from single nuclear spins can be filtered from much stronger background fields even at room temperature [6, 7, 8], thus representing a remarkable improvement compared to conventional NMR, which requires large nuclear ensembles and much stronger bias magnetic fields. This improvement comes at the expense that the probing system is now intrinsically non-classical, hence having knowledge of its time evolution is indispensable for the signal analysis. To overcome these challenges, an in-depth characterization of such ambiguous resonances experienced by the quantum sensor is necessary for the continued development of the field.
@@
Despite being applicable to two-level systems in general, DD techniques excel with color center defects in solids [14], given their versatility to sense different physical quantities and relative ease of experimental realization. Among them, the nitrogen vacancy (NV) center in diamond has been an epicenter of research due to its remarkable properties at a wide range of temperatures and compatibility with biological systems [15, 16], unlike for instance, superconducting quantum interference devices (SQUIDs) [17]. NV centers are a substitutional defects in the diamond lattice where a carbon atom is replaced by nitrogen 14N ($I^{n}=1$) or 15N ($I^{n}=1/2$) adjacent to a vacancy [Fig. 1 (b)]. The NV-, denoted as NV for simplicity, forms a spin triplet configuration ($S=1$), with well-defined quantum states within the bandgap of diamond [18]. They can be efficiently polarized to the $m_{s}=0$ state with a non-resonant green laser excitation via an optical pumping mechanism [19], which can also be used to read the population difference between $m_{s}=0$ and $m_{s}=\pm 1$ based on the resulting fluorescence. Finally, coherent transitions between the $m_{s}$ states can be achieved by applying resonant microwave (MW) pulses with controllable phase and duration [20]. This way, a DD sequence consists of three parts: laser initialization, MW manipulation and fluorescence readout. For NVs, $\pi$/2-pulses are included before and after the sequence, in order to drive the electronic spin to and from the plane perpendicular to the quantization axis.
@@
In the context of single NV multipulse sequences, ambiguous resonances can arise from different factors. We begin with Sec. II giving a general theoretical framework used in our simulations, following up with a description of the experimental and simulation methods in Sec. III. We then discuss the various ambiguous resonances appearing due to either external magnetic field misalignment and hyperfine interaction with 15N nuclei (Sec. IV.1), or coupling with lattice 13C having $I^{c}=1/2$ (Sec. IV.2), or free evolution of the system during finite pulses duration [13] (Sec. IV.3). In sight of the numerous possible ambiguous resonances specific to experimental parameters, we conclude with Sec. V providing a dataset with graphical interface where simulations can be visualized and compared with the user’s experimental data for spectral disambiguation of resonances. Although the discussion is focused on NVs, some of these effects are also relevant to other color centers and quantum systems in general.
@@ II Theoretical Framework
In contrast to conventional NMR, in DD quantum sensing, the probing system is an inseparable part from the target spin and the exact Hamiltonian of the quantum sensor needs to be taken into account. The total Hamiltonian of a NV in a multipulse sequence can be decomposed in three parts as $H=H_{0}+H_{1}(t)+H_{2}(t)$, where $H_{0}$ is the internal component, $H_{1}(t)$ is the control Hamiltonian describing the interaction with the MW pulses and $H_{2}(t)$ the sensing interaction itself. For the first internal component, we consider a spin model extensively studied and characterized [21, 22, 23]. It can be expressed in the Hilbert space of the electronic and nitrogen nuclear spins $\mathcal{H}_{S}\otimes\mathcal{H}_{I^{n}}$ in factors of decreasing energy as
$H_{0}=D\left({S_{z}}^{2}-\frac{S^{2}}{3}\right)$ (1.1)
$-\gamma^{e}\vec{B}_{0}\cdot\vec{S}$ (1.2)
@@
where the $z$ axis is taken along the NV axis. The first line (Eq. 1.1) corresponds to the zero-field splitting due to the dipolar interaction between the two electrons, where $D=2.87$ GHz is the axial component. The second and fourth lines (Eqs. 1.2 and 1.4) correspond to Zeeman interactions with the external magnetic field $\vec{B_{0}}$ for the electronic spin and nitrogen nuclear spin respectively, with gyromagnetic ratios $\gamma^{e}=-28025$ MHz/T, $\gamma^{n}=-4.316$ MHz/T for 15N and $\gamma^{n}=3.077$ MHz/T for 14N. Here we consider a general case where $\vec{B}_{0}$ is not necessarily aligned with the NV axis, as usually assumed. The third line (Eq. 1.3) is the hyperfine interaction between spins which is diagonal in this symmetry reference, with $a_{\perp}^{n}=-2.70$ MHz and $a_{\parallel}^{n}=-2.14$ MHz for 14N, while $a_{\perp}^{n}=3.65$ MHz and $a_{\parallel}^{n}=3.03$ MHz for 15N. When considering 14N, there is an extra quadrupole term given by $Q(I_{z}^{n})^{2}$, where $Q=-5.01$ MHz. Although small, the presence of the quadrupole interaction fundamentally changes the dynamics of the system, as the interaction tends to fix the nuclear spin orientation and suppress the Larmor precession of 14N nuclear spins for magnetic fields $B_{0}\ll Q/\gamma^{n}=1.63$ T [24].
The control Hamiltonian $H_{1}(t)$ for an oscillating field along a vector $\vec{B}_{1}$, with angular frequency $\Omega_{MW}$ and phase $\phi_{1}$ can be modeled as
@@
Typically, a rotating wave approximation (RWA) is assumed [25]. Thus, by setting the pulse frequency in resonance with one of the $m_{s}=0\leftrightarrow\pm 1$ transitions, $H_{1}(t)$ can be treated as a time independent operator which induces a rotation on the Bloch sphere of the respective manifold, where the pulse duration controls the angle and $\phi_{1}$ the axis of rotation in the rotating frame. Although this approximation often results in accurate descriptions of the electronic states transitions, it can overlook important interactions which lead to the ambiguous resonant responses. In particular, the $S_{x,y}$ components of the electronic Zeeman and the perpendicular terms of the hyperfine interactions $a_{\perp}^{n}\left(S_{x}I_{x}^{n}+S_{y}I_{y}^{n}\right)$ do not commute with the rotation operator and result in extra resonances (Sec. IV.1). This time dependency resulting from the non-commuting terms could still be handled as a perturbation to the secular part of the Hamiltonian up to 2nd order of the dominant term $(1/D)^{p}$ [24, 26, 27]. However, this approach is also limited, as it overlooks the possibility of double quantum transitions caused by multi-frequency or broad band excitations [28], as well as it implies small bias fields $|\gamma_{e}\vec{B}_{0}|\ll D$ and weak driving regimes $\gamma_{e}B_{1}\ll D$, which is not always experimentally desired due to extended pulse durations [29]. Furthermore, perturbation theory, in such circumstances, restricts on calculating approximated eigenvalues for the perturbed Hamiltonian, ignoring exact characteristics of the MW field and the free-evolution of the system during finite pulse durations. Which also leads to other ambiguous resonances (Sec. IV.3) [13].
@@
For a single nuclear spin in the strong coupling regime but without contact term [9], the sensing Hamiltonian $H_{2}$ can be simply expressed in an extended Hilbert space to include the sensed target spin $\mathcal{H}_{S}\otimes\mathcal{H}_{I^{n}}\otimes\mathcal{H}_{I^{c}}$, where we denote the sensed spin operator $I^{c}$ as carbon nuclei for a concrete example. For a spin-1/2 and neglecting the interaction between the nitrogen and target nuclear spin, this Hamiltonian is given by a hyperfine coupling between the electronic spin operator and the sensed nuclear spin added to the Zeeman interaction with the external magnetic field
$H_{2}=\vec{S}\cdot A^{c}\cdot\vec{I}^{c}-\gamma^{c}\vec{B}_{0}\cdot\vec{I}^{c}.$ (3)
@@
Diagonalizing the total Hamiltonian $H_{0}+H_{2}$ results in distinct target nuclear spin Larmor frequencies $\tilde{\omega}_{m_{s}}^{c}$, dependent on the electronic spin state. Given that the hyperfine matrix $A^{c}$ must be symmetrical, the sensing interaction is completely described by its six components plus the gyromagnetic ratio $\gamma^{c}$. The same Hamiltonian can also be used to describe an ensemble of nuclear spins, provided that their states are represented by density operators. Apart from nuclear spins, it is also desirable to sense oscillating classical fields. Particularly for this study, an externally applied RF signal has the advantage of having controllable intensity and frequency. In this case, the interaction is described in terms of a field $\vec{B}_{2}(t)$ with frequency $f_{0}$ as
@@
In a simplified manner, an ensemble of nuclear spins could also be thought of in terms of a semi-classical oscillating field [9, 30]. Despite their average magnetization being close to zero at room temperature, their collective precession around $\vec{B}_{0}$ results in an effective oscillating magnetic field with statically varying amplitude and phase, but with frequencies centered around the effective nuclear Larmor frequencies $\tilde{\omega}_{m_{s}}^{c}$ at each electronic state $m_{s}$. On the other hand, a weak classical RF signal could emulate a precessing nuclear spin and be approximately described in terms of an effective hyperfine interaction. Hence, an analogy can be drawn between the quantum and semi-classical descriptions from Eqs. 3 and 4, where the interaction is either represented in terms of the six components of the 3-dimensional symmetric hyperfine coupling matrix $A^{c}_{ij}$ plus the gyromagnetic ration $\gamma_{s}$, or the three nuclear precession frequencies $\tilde{\omega_{0,\pm 1}^{c}}$ plus four parameters of the effective field vector $\phi_{2}$ and $\vec{B}_{0}(x,y,z)$. Strictly speaking, both Hamiltonians describe fundamentally different processes, yet they share mathematical similarities and can lead to resembling effects at low order XY8-$M$ sequences, as further shown in Appendix E.